# Revealing Nanostructures in High-Entropy Alloys via Machine-Learning Accelerated Scalable Monte Carlo Simulation

### v2

@@ Abstract
The computational cost of traditional first-principles method quickly becomes prohibitively expensive as the number of atoms increases. This challenge is further amplified by the need to evaluate finite-temperature properties with Monte Carlo (MC) simulations, which is inherently challenging to parallelize due to sequential Markov chain updates. Here, we introduce Scalable Monte Carlo (SMC), an efficient MC simulation method that overcomes the parallelization bottlenecks in conventional MC simulation, reducing the computational complexity of a MC sweep from quadratic to linear. We present a GPU implementation of the SMC method, SMC-GPU, which simultaneously harnesses the thousands of processing cores on a GPU to accelerate the computation. By adopting a data-driven workflow that surrogates the computationally expensive density functional theory (DFT) with ML models, we demonstrate that SMC-GPU is capable of simulating systems of more than one-billion atoms, while maintaining the accuracy of first-principles methods. Using this unprecedented capability, we performed billion-atom MC simulations to investigate the nanostructure evolution of two important high-entropy alloys (HEAs), FeCoNiAlTi and MoNbTaW, in which the NPs are believed to be responsible for their superb mechanical properties. Our results reveal a rich diversity of nanostructures, including nanoparticles (NP), 3D-connected NP, and disorder protected nanophases. We quantitatively analyze the size, composition, and morphology of the nanostructures, as well as directly simulate the atom-probe-tomography (APT) needle. The results align well with available experimental observations. This work underscores the promising potential of leveraging large-scale MC simulation to explore the largely uncharted territory of nanostructure evolution in HEAs.
@@ 1 Introduction
One common strategy to accelerate atomistic simulations is to replace computationally expensive electronic structure calculations with efficient atomistic models. From the perspective of renormalization group theory [4, 5], the atomistic models can be interpreted as the effective model obtained by integrating out the electronic degrees of freedom, which substantially reduces the computational cost. Indeed, traditional empirical models, such as the embedded atom model (EAM) and cluster expansion methods, have been widely employed for large-scale thermodynamic simulations. However, traditional empirical models are limited in terms of functional form, which restricts their generalization capability and predictive accuracy. On the other hand, the advent of machine learning (ML) for atomistic systems [6, 7, 8, 9] has introduced a new paradigm to the construction of effective atomistic models [10, 11, 12, 13, 14]. Instead of being predefined, the parameters in ML atomistic models are automatically determined by training from high-quality first-principles datasets, therefore rendering it possible for the model to automatically capture complex interatomic interactions [15, 16, 17, 18]. Well-trained ML models are typically orders of magnitude faster than DFT methods, while can still retain their high accuracy [19, 20, 21]. This renders it possible to carry out high-accuracy atomistic simulations using supercells containing millions of atoms [22, 23, 24, 25], far exceeding the thousands-of-atoms limits of conventional DFT approaches.
@@
However, most machine learning potentials have been developed primarily to accelerate molecular dynamics (MD) simulations [22, 23, 7, 26, 27], with relatively limited applications in Monte Carlo (MC) simulations [28, 29, 16, 30, 29]. This is notable given that MC represents one of the two cornerstone methods in atomistic simulations, alongside MD [31]. A key challenge lies in the intrinsic sequential updating nature of widely used MC algorithms, such as the Metropolis algorithm. The MC trials are generally attempted site by site, which hinders large-scale parallelization. This is in stark contrast to MD, where all the atoms are updated simultaneously in a single step. The sequential updating nature of MC quickly becomes the bottleneck as the system size grows, which undermines the efficiency advantage of ML atomistic models when integrating the two. On the other hand, for many interesting phenomena [32, 33, 34], such as order-disorder transitions in chemically complex materials, MC simulations remain the only viable approach due to their sampling efficiency. Consequently, developing highly scalable MC algorithms that overcome the limitations of sequential updating is of paramount importance for realizing the potential of ML atomistic models.
@@
To overcome the parallelization bottleneck inherent in conventional MC simulations, we introduce the Scalable Monte Carlo (SMC) method, which is a generalized checkerboard algorithm [35] that reduces the computational complexity of an MC sweep from $O(N^{2})$ to $O(N)$. Compared to the checkerboard algorithm variant [35] that only applies to Ising interactions, we demonstrate that the SMC method can be generalized to arbitrary local interaction models by introducing a local interaction zone (LIZ), which makes it suitable for integration with machine learning atomistic models. We employ our method to study high-entropy alloys (HEAs) [36, 37], a class of chemically complex materials that have received significant attention due to their exceptional mechanical properties. These properties include overcoming the traditional strength-ductility trade-off [38, 39, 40], attributed to phenomena such as chemical short-range order [41, 42, 43], nanoprecipitates [39], and nanophases [44]. Understanding these features necessitates large-scale MC simulations. Beyond mechanical properties, the size and morphology of nanostructures in HEAs also offer promising opportunities for catalysis, sparking widespread interest in recent years [45, 46, 47]. We present a GPU-accelerated implementation of the SMC method, SMC-GPU, and apply it to investigate the nanostructure evolution in the $\rm{Fe_{29}Co_{29}Ni_{28}Al_{7}Ti_{7}}$ and MoNbTaW HEAs, employing two simple machine learning energy models using the local short-range order parameters [48, 33, 49] as input features. We demonstrate the excellent efficiency and scalability of the SMC method, which enables the simulation of atomistic systems exceeding one billion atoms using a single GPU (graphic processing unit). Our results reveal a rich diversity of interesting nanoscale phenomena, including nanoparticles (NP), 3D-connected NP network, and disorder protected nanophases. We quantitatively analyze the size, composition, and morphology of the nanostructures, which align well with available experimental results obtained with atom-probe-tomography (APT) and electron microscopy. Finally, our results reveal that the intricate nanoscale interplay of order and disorder in high-entropy alloys (HEAs) stems from the combined effects of chemical complexity and temperature, offering valuable guidance for alloy design.
@@ 2.1 Performance of the SMC algorithm
The SMC algorithm can be seen as a generalization of the checkerboard algorithm for the 2D Ising model, as explained in Method 3.2 and Supplementary 5.1. The key insight lies in rendering sequential Monte Carlo (MC) trials independent by partitioning the system into sufficiently large link-cells, thereby isolating the influence of individual MC trial moves. To simplify the explanation, a schematic of the link-cell for the 2D square lattice is shown in Fig. 1 (a). Each site is labeled by two indices $(i_{C},i_{A})$, where $i_{C}$ represents different link-cells and $i_{A}$ represents the different atoms within the link-cell. For the sake of discussion, let us assume that the atoms have only nearest-neighbor interactions. In such a case, the MC updates of the atoms with the same $i_{A}$ index but different $i_{C}$ indexes are independent of each other. This obviously presents a parallelism opportunity. For instance, consider atom $(1,13)$. During MC update, atom $(1,13)$ can swap with any of its nearest neighbors $(1,12)$, $(1,14)$, $(1,8)$ and $(1,18)$, which are colored as yellow in Fig. 1(a). Due to the short interaction range (nearest-neighbor), these moves would have no impact on calculating the energy changes for the swap moves of atom $(i_{C},13)$ (red sites in Fig. 1(a)), and the degree of parallelism is the number of cells $n_{C}$. We denote the yellow and green sites surrounding each red site in Fig. 1 as a local-interaction zone (LIZ), which is a name inspired by the locally self-consistent multiple scattering (LSMS) method [50, 51]. In other words, for any two sites, as long as the MC trial of one site does not affect the energies of the LIZ of the other site, then these two sites are independent. It is easy to see that the above discussion based on the 2D square lattice can be extended to the case of 3D crystal. Furthermore, in the SMC method, a domain decomposition scheme is introduced to distribute the lattice among multiple GPUs, which further enhances the achievable system size, as illustrated in Fig. 1 (b). Generally, to determine the energy change resulting from the MC swap trial at site $i$, the local energies of each site within the local interaction zone of site $i$ must be evaluated. Additionally, chemical environment information extending beyond the LIZ is required for the energy change calculation, as illustrated by the LIZ+ region in Fig. 1 (c). Note that for a pair-interaction model, the above discussion can be simplified since the total energy changes can be calculated from the local energies of the two sites involved in the swap trial. As a result, the speed of the effective pair-interaction (EPI) model is faster than the generalized nonlinear model by approximately a factor of 50, as shown in Fig. 1 (d).
@@
Figure 1: Schematics of the scalable Monte Carlo (SMC) method. (a) Illustrate the SMC method with a 2D square lattice, in which each site is denoted by two indices $(i_{C},i_{A})$, where $i_{C}$ represents different link-cells, and $i_{A}$ represents the different atoms within the $n_{A}$ atom cell. (b) The whole chemical configuration is distributed on multiple GPUs, and the atoms near the boundary need to be communicated between GPUs $n_{A}$ times in every MC sweep. (c) A 2D view of the fcc or bcc lattice, site 1 (red) can swap with each of its nearest neighboring sites (yellow), and the green sites represent the ones that the local energies can be affected. The yellow and green sites, together with the centering atom form the LIZ. The chemical environment with LIZ+ are needed to calculate the energy change due to a swap trial. (d) A log-log plot of computation time vs the system size to illustrate the speed-up ratio of the SMC method as compared to DFT (MuST-KKR), linear-DFT (LSMS), and GNN (Allegro). The measured values for SMC are signified as stars and ideal scaling is assumed for all lines.
@@
To illustrate the strength of our method, we compare the theoretical speed of the SMC method with other schemes to integrate MC with high-accuracy energy prediction methods, such as DFT, linear-DFT, and GNN (graph neural networks), as shown in Fig. 1 (d). Note that the lines are drawn by extrapolating from a data point using ideal scaling behavior. The actual computational speeds of the various methods are inherently influenced by many factors, such as hardware, software implementations, material systems, and computational parameters. Therefore, the discussion presented here is intended solely to provide a theoretical estimation of their relative performance in terms of orders of magnitude. The stars in the figure represent the measured values of the SMC method, and the lines show the ideal scaling behaviors, which are explained in Method 3.1. From Fig. 1 (d), we see that for a one million atom system, we have the speedup $S=10^{25}$, as highlighted in Fig. 1 (d). We can also see that for a nonlinear model, such as the Local_SRO model introduced in the following section, the speedup will be reduced by a factor of 50, due to the aforementioned necessity of evaluating the local energies of each site within the LIZ. If we increase the system size to one billion atoms, then the speedup of SMC with respect to DFT will be further increased by a factor of $10^{9}$ to reach $10^{34}$, as highlighted in Fig. 1 (d). This analysis clearly demonstrates the exceptional performance of the SMC method for large-scale MC simulations.
@@
In addition to theoretical estimations, we compare our work with selected previous studies, as summarized in Tab. 1. It can be seen that our work represents the largest system size achieved in atomistic simulations at ab initio accuracy. Such capability enables us to directly study large nanostructures comprising of millions of atoms, in contrast to previous works that are limited to short-range order (SRO), long-range order (LRO), and small nanoprecipitates. To contextualize the simulation scale of our work, we extend our search beyond HEAs to identify the largest atomistic simulation system reported in the literature. A recent study achieved this with a system of 29 billion atoms, which employs 35 million CPU cores on one of the world’s largest supercomputers (comprising 90,000 computing nodes, approximately 84% of the entire system). In contrast, constrained by computational resources, our work utilized only two NVIDIA H800 GPUs, yet already achieved a simulation scale of one billion atoms. Note that this already surpasses the 100-million-atom system size of the 2020 Gordon-Bell Prize winner [22]. Furthermore, the comparison presented above demonstrates that, through careful design of the Monte Carlo algorithm to fully exploit its inherent parallelization potential, MC simulations can rival molecular dynamics (MD) methods in terms of scalability, which are traditionally viewed as advantageous over MC methods in simulation scale. Finally, we highlight that the excellent computational efficiency of the SMC method not only presents a success in pushing the simulation length scale forward, but also provides a particularly important tool for understanding the vital role of nanostructures [52, 40] in the exceptional mechanical properties in HEAs [47, 53, 39], such as overcoming the trade-off in strength and ductility [52, 39].

### v3

@@ Abstract
First-principles Monte Carlo (MC) simulations at finite temperatures are computationally prohibitive for large systems due to the high cost of quantum calculations and poor parallelizability of sequential Markov chains in MC algorithms. We introduce scalable Monte Carlo at eXtreme (SMC-X), a generalized checkerboard algorithm designed to accelerate MC simulation with arbitrary short-range interactions, including machine learning potentials, on modern accelerator hardware. The GPU implementation, SMC-GPU, harnesses massive parallelism to enable billion-atom simulations when combined with machine-learning surrogates of density functional theory(DFT). We apply SMC-GPU to explore nanostructure evolution in two high-entropy alloys, FeCoNiAlTi and MoNbTaW, revealing diverse morphologies including nanoparticles, 3D-connected NPs, and disorder-stabilized phases. We quantify their size, composition,and morphology, and simulate an atom-probe tomography (APT) specimen for direct comparison with experiments. Our results highlight the potential of large-scale, data-driven MC simulations in exploring nanostructure evolution in complex materials, opening new avenues for computationally guided alloy design.
@@ 1 Introduction
One common strategy to accelerate atomistic simulations is to replace computationally expensive electronic structure calculations with efficient atomistic models. From the perspective of renormalization group theory [4, 5, 6], the atomistic models can be interpreted as the effective model obtained by integrating out the electronic degrees of freedom, which substantially reduces the computational cost. Indeed, traditional empirical models, such as the embedded atom model (EAM) and cluster expansion methods, have been widely employed for large-scale thermodynamic simulations. However, traditional empirical models are limited in terms of functional form, which restricts their generalization capability and predictive accuracy. On the other hand, the advent of machine learning (ML) for atomistic systems [7, 8, 9, 10] has introduced a new paradigm to the construction of effective atomistic models [11, 12, 13, 14, 15]. Instead of being predefined, the parameters in ML atomistic models are automatically determined by training from high-quality first-principles datasets, therefore rendering it possible for the model to automatically capture complex interatomic interactions [16, 17, 18, 19]. Well-trained ML models are typically orders of magnitude faster than DFT methods, while can still retain their high accuracy [20, 21, 10]. This renders it possible to carry out high-accuracy atomistic simulations using supercells containing millions of atoms [22, 23, 24, 25], far exceeding the thousands-of-atoms limits of conventional DFT approaches.
@@
However, most machine learning potentials have been developed primarily to accelerate molecular dynamics (MD) simulations [22, 23, 8, 9, 26, 27], with relatively limited applications in Monte Carlo (MC) simulations [28, 29, 17, 30, 29]. This is notable given that MC represents one of the two cornerstone methods in atomistic simulations, alongside MD [31]. A key challenge lies in the intrinsic sequential updating nature of widely used MC algorithms, such as the Metropolis algorithm. The MC trials are generally attempted site by site, which hinders large-scale parallelization. This is in stark contrast to MD, where all the atoms are updated simultaneously in a single step. The sequential updating nature of MC quickly becomes the bottleneck as the system size grows, which undermines the efficiency advantage of ML atomistic models when integrating the two. On the other hand, for many interesting phenomena [32, 33, 34], such as order-disorder transitions in chemically complex materials, MC simulations remain the only viable approach due to their sampling efficiency. Consequently, developing highly scalable MC algorithms that overcome the limitations of sequential updating is of paramount importance for realizing the potential of ML atomistic models.
Model Model Type Chip Distributed Year Journal & Ref.
Lennard-Jones Pairwise CPU Y 1986 Thesis [35]
Lennard-Jones Pairwise CPU Y 1996 JCC [36]
Ising NN GPU N 2009 JCP [37]
Ising NN GPU Y 2010 CPC [38]
Cluster-expasnion Short-range CPU Y 2012 PRB [39]
Lennard-Jones Pairwise GPU N 2013 JCP [40]
Ising NN FPGA N 2013 JCP[41]
Hard-disk Hard core GPU N 2013 JCP [42]
Hard-disk Hard core GPU Y 2016 CPC [43]
Coulomb Long-range GPU N 2017 JCP[44]
Ising NN TPU Y 2019 SC [45]
Ising NN GPU-TC Y 2020 CPC [46]
Magnetic Long-range GPU Y 2023 JAP [47]
EPI, qSRO Short-range GPU Y 2025 This work
Table 1: Comparison of the method proposed in this work with previous checkerboard variants of parallel MC simulation, for atomistic simulation of materials. Algorithms generalizable to MLP, implemented on accelerators, or distributed over multiple chips, are highlighted with bold font. In the model type column, NN refers to nearest-neighbor.
A widely adopted technique to address the parallelization challenge in Monte Carlo (MC) simulations is the checkerboard algorithm, originally developed for the Ising model with nearest-neighbor interactions. The combination of Ising model and checkerboard algorithm presents great parallel opportunities at core level, and has been successfully applied across various accelerators, including GPUs [37, 38], GPU tensor cores [46], TPUs [45], and ASICs [41]. Extending the checkerboard algorithm to non-Ising models poses greater challenges for achieving efficient parallelism. One of the earliest such efforts was by M. Johnson [35] for Lennard-Jones potential. Later, J. Anderson and collaborators proposed an extension for off-lattice hard-disk systems [42, 43]. However, a key limitation of these efforts is their restriction to pairwise interactions, making them unsuitable for machine-learned potentials (MLPs), which often involve many-body interactions. For general short-range interactions, B. Sadigh and co-authors introduced the Scalable Parallel Monte Carlo (SPMC) algorithm [39], which supports higher-order interactions such as those in cluster expansion models. Their implementation was later integrated into the hybrid MC-MD framework in the LAMMPS package [48] and has been widely adopted for use with MLPs, including models like SNAP [49] and MTP [50]. Nevertheless, the original SPMC implementation was designed for CPU architectures, not GPUs. Moreover, it relies on a static domain decomposition scheme, which limits the parallel scalability. Finally, the practice of confining the local trial moves to non-interacting domains produces weak spatial correlations that can slow down the approach to equilibrium [39].
To overcome the above challenges, we introduce the Scalable Monte Carlo at eXtreme (SMC-X) method. Compared to previous methods summarized in Tab. 1, SMC-X is unique in that it is designed for both two targets: (1) efficient implementation on accelerators, (2) support for arbitrary short-range interactions, such as those found in machine-learned potentials (MLPs). The key innovation in SMC-X is the introduction of the Local Interaction Zone (LIZ) within the link-cell (LC) strategy (a checkerboard variant), which enables efficient execution on accelerators. Although SMC-X shares conceptual similarities with the SPMC method [39], it avoids the key limitations of that approach mentioned above. In SMC-X, the LIZ+LC region is dynamically centered on the atoms updated in each mini-sweep, which brings two major advantages: (1) The parallelism scales linearly with the number of atoms, far exceeding the core-limited scalability of SPMC, and making it well-suited for high-throughput accelerators such as GPUs; (2) All sites are treated uniformly, maintaining equivalence with serial Monte Carlo updates (aside from update order, which does not affect detailed balance) and avoiding the spatial correlation issues inherent in SPMC.
@@
To demonstrate the capability of SMC-X in tackling challenging materials science problems previously intractable due to limitations in spatial and temporal scales, we employ SMC-X to study the nanoparticles in high-entropy alloys (HEAs) [51, 52]. HEA is a class of chemically complex materials that have received significant attention due to their exceptional mechanical properties. These properties include overcoming the traditional strength-ductility trade-off [53, 54, 55], attributed to phenomena such as chemical short-range order [56, 34, 57], nanoprecipitates [54], and nanophases [58]. Understanding these features necessitates large-scale MC simulations. Beyond mechanical properties, the size and morphology of nanostructures in HEAs also offer promising opportunities for catalysis, sparking widespread interest in recent years [59, 60, 61]. As a general algorithm, SMC-X can be applied to different accelerators. In this work, we tailored the parallelization strategies in SMC-X for efficient execution on GPUs, and refer to this implementation as SMC-GPU. Specifically, we apply SMC-GPU to investigate the nanostructure evolution in the $\rm{Fe_{29}Co_{29}Ni_{28}Al_{7}Ti_{7}}$ and MoNbTaW HEAs, employing two simple machine learning energy models using the local short-range order parameters [62, 33, 63] as input features. We demonstrate the excellent efficiency and scalability of the SMC-X method, which enables the simulation of atomistic systems exceeding one billion atoms using a single GPU (graphic processing unit). Our results reveal a rich diversity of interesting nanoscale phenomena, including nanoparticle (NP), 3D-connected NPs, and disorder protected nanophases. We quantitatively analyze the size, composition, and morphology of the nanostructures, which align well with available experimental results obtained with atom-probe-tomography (APT) and electron microscopy. Finally, our results reveal that the intricate nanoscale interplay of order and disorder in high-entropy alloys (HEAs) stems from the combined effects of chemical complexity and temperature, offering valuable guidance for alloy design.
@@ 2.1 Performance of the SMC-X algorithm
The SMC-X algorithm can be seen as a generalization of the checkerboard algorithm for the 2D Ising model, as explained in Method 3.2 and Supplementary 5.1. The key insight lies in rendering sequential Monte Carlo (MC) trials independent by partitioning the system into sufficiently large link-cells, thereby isolating the influence of individual MC trial moves. To simplify the explanation, a schematic of the link-cell for the 2D square lattice is shown in Fig. 1 (a). Each site is labeled by two indices $(i_{C},i_{A})$, where $i_{C}$ represents different link-cells and $i_{A}$ represents the different atoms within the link-cell. For the sake of discussion, let us assume that the atoms have only nearest-neighbor interactions. In such a case, the MC updates of the atoms with the same $i_{A}$ index but different $i_{C}$ indexes are independent of each other. This obviously presents a parallelism opportunity. For instance, consider atom $(1,13)$. During MC update, atom $(1,13)$ can swap with any of its nearest neighbors $(1,12)$, $(1,14)$, $(1,8)$ and $(1,18)$, which are colored as yellow in Fig. 1(a). Due to the short interaction range (nearest-neighbor), these moves would have no impact on calculating the energy changes for the swap moves of atom $(i_{C},13)$ (red sites in Fig. 1(a)), and the degree of parallelism is the number of cells $n_{C}$. We denote the yellow and green sites surrounding each red site in Fig. 1 as a local-interaction zone (LIZ), which is a name inspired by the locally self-consistent multiple scattering (LSMS) method [64, 65]. In other words, for any two sites, as long as the MC trial of one site does not affect the energies of the LIZ of the other site, then these two sites are independent. It is easy to see that the above discussion based on the 2D square lattice can be extended to the case of 3D crystal. Furthermore, in the SMC-X method, a domain decomposition scheme is introduced to distribute the lattice among multiple GPUs, which further enhances the achievable system size, as illustrated in Fig. 1 (b). Generally, to determine the energy change resulting from the MC swap trial at site $i$, the local energies of each site within the local interaction zone of site $i$ must be evaluated. Additionally, chemical environment information extending beyond the LIZ is required for the energy change calculation, as illustrated by the LIZ+ region in Fig. 1 (c). Note that for a pair-interaction model, the above discussion can be simplified since the total energy changes can be calculated from the local energies of the two sites involved in the swap trial. As a result, the speed of the effective pair-interaction (EPI) model is faster than the generalized nonlinear model by approximately a factor of 50, as shown in Fig. 1 (d).
@@
Figure 1: Schematics of the scalable Monte Carlo (SMC-X) method. (a) Illustrate the SMC-X method with a 2D square lattice, in which each site is denoted by two indices $(i_{C},i_{A})$, where $i_{C}$ represents different link-cells, and $i_{A}$ represents the different atoms within the $n_{A}$ atom cell. (b) The whole chemical configuration is distributed on multiple GPUs, and the atoms near the boundary need to be communicated between GPUs $n_{A}$ times in every MC sweep. (c) A 2D view of the fcc or bcc lattice, site 1 (red) can swap with each of its nearest neighboring sites (yellow), and the green sites represent the ones that the local energies can be affected. The yellow and green sites, together with the centering atom form the LIZ. The chemical environment with LIZ+ are needed to calculate the energy change due to a swap trial. (d) A log-log plot of computation time vs the system size to illustrate the speed-up ratio of the SMC-X method as compared to DFT (MuST-KKR), linear-DFT (LSMS), and GNN (Allegro). The measured values for SMC-X are signified as stars and ideal scaling is assumed for all lines.
@@
To illustrate the strength of our method, we compare the theoretical speed of the SMC-X method with other schemes to integrate MC with high-accuracy energy prediction methods, such as DFT, linear-DFT, and GNN (graph neural networks), as shown in Fig. 1 (d). Note that the lines are drawn by extrapolating from a data point using ideal scaling behavior. The actual computational speeds of the various methods are inherently influenced by many factors, such as hardware, software implementations, material systems, and computational parameters. Therefore, the discussion presented here is intended solely to provide a theoretical estimation of their relative performance in terms of orders of magnitude. The stars in the figure represent the measured values of the SMC-X method, and the lines show the ideal scaling behaviors, which are explained in Method 3.1. From Fig. 1 (d), we see that for a one million atom system, we have the speedup $S=10^{25}$, as highlighted in Fig. 1 (d). We can also see that for a nonlinear model, such as the qSRO model introduced in the following section, the speedup will be reduced by a factor of 50, due to the aforementioned necessity of evaluating the local energies of each site within the LIZ. If we increase the system size to one billion atoms, then the speedup of SMC-X with respect to DFT will be further increased by a factor of $10^{9}$ to reach $10^{34}$, as highlighted in Fig. 1 (d). This analysis clearly demonstrates the exceptional performance of the SMC-X method for large-scale MC simulations.
@@
In addition to theoretical estimations, we compare our work with selected previous studies, as summarized in Tab. 2. It can be seen that our work represents the largest system size achieved in atomistic simulations at ab initio accuracy. Such capability enables us to directly study large nanostructures comprising of millions of atoms, in contrast to previous works that are limited to short-range order (SRO), long-range order (LRO), and small nanoprecipitates. To contextualize the simulation scale of our work, we extend our search beyond HEAs to identify the largest atomistic simulation system reported in the literature. A recent study achieved this with a system of 29 billion atoms, which employs 35 million CPU cores on one of the world’s largest supercomputers (comprising 90,000 computing nodes, approximately 84% of the entire system). In contrast, constrained by computational resources, our work utilized only two NVIDIA H800 GPUs, yet already achieved a simulation scale of one billion atoms. Note that this already surpasses the 100-million-atom system size of the 2020 Gordon-Bell Prize winner [22]. Furthermore, the comparison presented above demonstrates that, through careful design of the Monte Carlo algorithm to fully exploit its inherent parallelization potential, MC simulations can rival molecular dynamics (MD) methods in terms of scalability, which are traditionally viewed as advantageous over MC methods in simulation scale. Finally, we highlight that the excellent computational efficiency of the SMC-X method not only presents a success in pushing the simulation length scale forward, but also provides a particularly important tool for understanding the vital role of nanostructures [66, 55] in the exceptional mechanical properties in HEAs [61, 67, 54], such as overcoming the trade-off in strength and ductility [66, 54].